# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ 3.2.1 Frequency to wavenumber conversion
Under the assumption that the turbulent structures of different scales advect past the measurement location at similar speeds, one could essentially transform the frequencies ($f$) to wavenumbers ($\kappa$) using $\kappa=2\pi f/U$, where $U$ is the mean wind speed estimated from the sonic anemometer. Irrespective of whether one uses $U$ or $U_{C}$ to convert the frequencies to wavenumbers, it is a linear transformation. However, if the convective speeds of turbulent structures are scale-dependent, it introduces additional complexities to the transformation. From our DTS observations, it is clear that $U_{C}(\kappa z)$ is equal to $U_{C}$ at scales $\kappa z<1$, while it varies as ${(\kappa z)}^{-0.8}$ at scales $\kappa z>1$ (see Fig. 5a). One could express this dependence through the following functional form,
$\frac{U_{c}(\kappa z)}{U_{c}}=\frac{1}{{(1+\kappa z)}^{4/5}},$ (9)
which satisfies the aforementioned conditions. When compared against the observations, this functional form fits the data exceptionally well (see the yellow dash-dotted line in Fig. 5a). Now, if this relationship is used to convert the frequencies to wavenumbers, one could write,
@@
By replacing Eq. 9 in Eq. 10, and after some algebraic manipulations, one could rewrite Eq. 10 as,
@@
where $a=(2\pi fz/U_{C})^{5}$. By solving the polynomial in Eq. 11 and restricting its roots to positive values, one could essentially convert the frequencies to wavenumbers by using a scale-dependent convection speed.
@@
Before we discuss the impact of this transformation on turbulence spectra and cospectra, it is important to notice that since Eq. 11 is a non-linear function, the wavenumbers generated from Eq. 11 will not be spaced equally in a linear space. Accordingly, the wavenumber increments ($d\kappa$) themselves would depend on the values of $\kappa$. In fact, by differentiating Eq. 11, one can show,
$d\kappa=\frac{10\pi}{U_{C}}\frac{{(1+\kappa z)}^{9/5}}{5+\kappa z}df.$ (12)
Eq. 12 dictates that $d\kappa=f(\kappa)$ at a fixed $z$, meaning that constant $df$ does not convert to constant $d\kappa$. This creates an issue while estimating the spectral densities ($S_{xx}(\kappa)$) as those are defined as $S_{xx}(\kappa)=E_{xx}(\kappa)/d\kappa$, where $E_{xx}(\kappa)$ are the spectral powers of a turbulent signal $x$ and $d\kappa$ is the wavenumber increment independent of $\kappa$.
Therefore, to solve this paradox, we exponentially bin the scaled wavenumbers ($\kappa z$) generated from Eq. 11 and then compute the bin-averaged spectral powers ($E_{xx}(\kappa z_{\rm bin})$) corresponding to those wavenumber ranges, bounded between $\kappa z_{\rm bin}(i)<\kappa z<\kappa z_{\rm bin}(i)+d\log(\kappa z_{\rm bin})$, where $i$ spans from 1 to $N$ with $N$ being the number of exponential bins being used. Here, we use $N=50$ for our application. To get an equivalent estimate in the linear space, the spectral density at a wavenumber of $\kappa z_{\rm bin}(i)$ is expressed as, $S_{xx}(\kappa z_{\rm bin}(i))=E_{xx}(\kappa z_{\rm bin}(i))/[\kappa z_{\rm bin}(i)d(\log(\kappa z_{\rm bin})]$, where $\kappa z_{\rm bin}(i)$ itself appears in the denominator. This is necessary since $d(\log(\kappa))=d\kappa/\kappa$. The entire procedure described above is a standard one while dealing with empirical estimations of the probability density functions of stochastic variables that are known to display power-law variations (Chowdhuri et al., 2020a). Henceforth, while reporting the spectral results derived from the scale-dependent convection speeds, we omit the subscript from $\kappa z_{\rm bin}$.
@@ 3.2.2 DTS temperature spectra
Figure 6: The ensemble averaged temporal and spatial spectra of temperature fluctuations from the DTS measurements are shown for (a)–(e) along- and (f) perpendicular-wind cases. Subplots (a)–(e) show data from cable sections 1 to 5 and subplot (f) shows data from section 6. The spectral amplitudes ($S_{TT}$) are normalized with the variance of temperature ($\sigma^{2}_{T}$) computed from the EC measurements at the clearcut. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. For the temporal spectra, the frequency ($f$) to wavenumber ($\kappa$) conversion is carried out using bulk ($U_{C}$) and scale-dependent ($U_{c}(\kappa z)$) convection speeds, and shown as blue and black colored lines, respectively. For the spatial data, the wavenumbers are directly computed from the available spatial information (shown as red-colored lines). The Kolmogorov $-5/3$ prediction for the spectral slopes is shown for both temporal and spatial measurements.

### v2

@@ 3.2.1 Frequency to wavenumber conversion
Under the assumption that the turbulent structures of different scales advect past the measurement location at similar speeds, one could essentially transform the frequencies ($f$) to wavenumbers ($\kappa$) using $\kappa=2\pi f/U$, where $U$ is the mean wind speed estimated from the sonic anemometer. Irrespective of whether one uses $U$, $U_{C}$, or $U_{e}$ to convert the frequencies to wavenumbers, it is a linear transformation. However, if the convective speeds of turbulent structures are scale-dependent, it introduces additional complexities to the transformation. From our DTS observations, it is evident that $U_{C}(\kappa z)$ is equal to $U_{C}$ at scales $\kappa z<1$, while it varies as ${(\kappa z)}^{-0.8}$ at scales $\kappa z>1$ (see Fig. 5b). In general, one could express this dependence through the following functional form,
$\frac{U_{c}(\kappa z)}{U_{C}}=\frac{1}{{(1+\kappa z)}^{\alpha}},$ (7)
with $0\leq\alpha<1$. Here, $\alpha=1$ corresponds to a limit for the randomly-shuffled samples (see Fig. 11a) while $\alpha=0$ indicates no scale-dependency. If $\alpha=0.8$, the functional form of Eq. 7 fits the data exceptionally well (see the yellow dash-dotted line in Fig. 5b). Now, if this relationship is used to convert the frequencies to wavenumbers, one could write,
@@
By replacing Eq. 7 in Eq. 8, one could rewrite Eq. 8 as,
$\kappa z=\frac{2\pi fz}{U_{C}}{(1+\kappa z)}^{\alpha}.$ (9)
Depending on $\alpha$, Eq. 9 is a $\rm n$-th order polynomial of $\kappa z$. For instance, when $\alpha=0.8=4/5$, Eq. 9 becomes,
@@
where $a=(2\pi fz/U_{C})^{5}$. By solving the polynomial in Eq. 9 and restricting its roots to positive values, one could convert the frequencies to wavenumbers by using a scale-dependent convection speed.
@@
It is important to notice that since Eq. 9 is a non-linear function, the wavenumbers generated from Eq. 9 will not be spaced equally in a linear space. Accordingly, the wavenumber increments ($d\kappa$) themselves would depend on the values of $\kappa$. In fact, by differentiating Eq. 9, one can show,
$d(\kappa z)=\frac{{(1+\kappa z)}^{\alpha+1}}{1+(1-\alpha)\kappa z}d(\frac{2\pi fz}{U_{C}}).$ (11)
We next discuss the implications of Eqs. 9 and 11 on spectral conversion.
3.2.2 Spectral conversion
The spectral energy at a specific wavenumber $\kappa z$ ($E_{xx}(\kappa z)$, where $x$ could be either velocity or scalar) is evaluated by integrating the spectral energy density between the limits $\kappa z$ and $\kappa z+d(\kappa z)$, and expressed as (within a specific constant),
$E_{xx}(\kappa z)\sim S_{xx}(\kappa z)d(\kappa z),$ (12)
where $S_{xx}(\kappa z)$ is the spectral energy density. In the absence of spatial information, one could apply Taylor’s hypothesis to convert the frequencies to wavenumbers. In that case, one could write Eq. 12 as,
$E_{xx}(\kappa z)\sim S_{xx}(2\pi fz/U_{c})d(2\pi fz/U_{c}).$ (13)
Let us consider that in the inertial subrange scales ($\kappa z\gg 1$) the spectral energy density in the temporal domain vary as $f^{-\beta}$ while at larger scales ($\kappa z\ll 1$) it varies as $f^{-\gamma}$. Therefore, under such conditions, one could write Eq. 13 as,
$E_{xx}(\kappa z)\sim\begin{cases}{(2\pi fz/U_{c})}^{-\gamma}d(2\pi fz/U_{c})&{\kappa z\ll 1}\\ {(2\pi fz/U_{c})}^{-\beta}d(2\pi fz/U_{c})&{\kappa z\gg 1}\end{cases}$ (14)
The Taylor’s hypothesis is valid when $\alpha=0$. However, when $\alpha$ is non-zero, one needs to replace $2\pi fz/U_{c}$ and $d(2\pi fz/U_{c})$ with Eqs. 9 and 11, respectively. By doing such replacements, one ends up with,
$\begin{split}E_{xx}(\kappa z)&\sim{\left[\frac{\kappa z}{{(1+\kappa z)}^{\alpha}}\right]}^{-\gamma}\left[\frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}\right]{d(\kappa z)}\\ E_{xx}(\kappa z)&\sim{\left[\frac{\kappa z}{(\kappa z)^{\alpha}(1+\frac{1}{\kappa z})^{\alpha}}\right]}^{-\beta}\left[\frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}\right]{d(\kappa z)}\end{split}$ (15)
When $\kappa z\ll 1$,
$\begin{split}{(1+\kappa z)}^{\alpha}&\to 1\\ \frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}&\to 1.\end{split}$ (16)
However, at inertial subrange scales, $\kappa z\gg 1$, and hence one could simplify the terms in Eq. 15 as,
$\begin{split}\left(1+\frac{1}{\kappa z}\right)^{\alpha}&\to 1\\ {(1+\kappa z)}^{\alpha+1}&={(\kappa z)}^{\alpha+1}\left(1+\frac{1}{\kappa z}\right)^{\alpha+1}\to{(\kappa z)}^{\alpha+1}\\ [1+(1-\alpha)\kappa z]&\to(1-\alpha)\kappa z\end{split}$ (17)
By inserting these approximations of Eqs. 16 and 17 in Eq. 15, one obtains,
$E_{xx}(\kappa z)\sim\begin{cases}{(\kappa z)}^{-\gamma}d(\kappa z)&{\kappa z\ll 1}\\ (1-\alpha){(\kappa z)}^{-\beta(1-\alpha)}(\kappa z)^{-\alpha}d(\kappa z)&{\kappa z\gg 1}\end{cases}$ (18)
For $\kappa z\gg 1$, Eq. 18 could be further rearranged as,
$E_{xx}(\kappa z)\sim(1-\alpha){(\kappa z)}^{-\beta+(\beta-1)\alpha}d(\kappa z).$ (19)
This is the desired result and one could see that the scale-dependent conversion does not change the spectral slope at smaller wavenumbers ($\kappa z\ll 1$) but yields a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange ($\kappa z\gg 1$) when the temporal frequencies are converted to wavenumbers using Eq. 7. The spectral behaviors described by Eq. 18 are valid for low- and high-wavenumber ranges. For the mid-wavenumber range, additional terms will appear after carrying out the binomial expansions of ${(1+\kappa z)}^{\alpha}$ and ${(1+\kappa z)}^{\alpha+1}$. Specific to the inertial subrange ($\kappa z\gg 1$), if $\beta>1$, the extra factor of $(\beta-1)\alpha$ in Eq. 19 implies that applying Taylor’s hypothesis underestimates the turbulent spectrum in the inertial subrange when $\alpha$ is non-zero. By comparing Eq. 19 with the findings of cheng2017failure, they obtained $\beta=2$ and $\alpha=1/3$ in the inertial subrange scales (see their Figs. 3 and 5). After replacing these values in Eq. 19, one obtains $E_{xx}(\kappa z)\sim{(\kappa z)}^{-5/3}d(\kappa z)$, which is precisely the result of cheng2017failure (see Section 2 in the Supplementary Material where we establish this equivalence). For our flow set up, $\beta$ appears to be $5/3$ (see Fig. 12a), and therefore the conversion yields, $5/3+2\alpha/3$. Note that a similar conversion can be carried out for the flux cospectra, with only difference being the slope $\beta$ will have a different value than the spectra (wyngaard1972cospectral). An important aspect to note is for the values of $0\leq\alpha<1$ and $\beta\geq 1$, the upper bound of the spectral slope in Eq. 19 is always 1-1. In essence, by applying Eqs. 9 and 11 on any temporal turbulence dataset, one could convert the spectra from temporal to spatial domain at any specific value of $\alpha$.
@@ 3.2.3 DTS temperature spectra
Figure 6: Corresponding to section-1 of the DTS cable, (a) the original and converted DTS temporal spectra are displayed. The original DTS temporal spectrum is shown as blue lines with square markers, while the converted one using scale-dependent convection speeds is shown as a black line with square markers. The conversion yields a spectral slope of $5/3+0.8\times 2/3\approx-1.13$ in the temporal domain (see Eq. 19), illustrated as a magenta line. (b) The spatial DTS spectrum is shown, and the magenta line there indicates the same slope as obtained in the temporal domain after conversion.